# Reasoning on Time-Series for Financial Technical Analysis

### v1

@@ 6. Conclusion
In this work, we tackled the task of verbally reasoning over financial time-series data. This task is challenging as it switches between the time-series and natural language domain for the stock price data and the reasoning step. To deal with this, we introduce our Verbal Technical Analysis (VTA) framework, which combines verbal and latent reasoning to produce interpretable time-series forecasts. The framework utilizes our Time-GRPO method to finetune the reasoning model, and conditions its forecasts on the reasoning attributes. We conducted extensive experiments across multiple benchmarks and find VTA can achieve state-of-the-art forecasting accuracy while also producing high-quality reasoning traces.
Appendix A List of Technical Indicators
Table 6. A list of the financial technical indicators used in Time-GRPO. Indicator Description Formula Simple Moving Average (SMA) Identifies trend by smoothing price data over a period $\text{SMA}=\frac{1}{n}\sum_{i=1}^{n}\text{Price}_{i}$ Exponential Moving Average (EMA) Measures the trend by smoothing price data with greater weight to recent prices $\text{EMA}_{t}=\text{Price}_{t}\cdot\alpha+\text{EMA}_{t-1}\cdot(1-\alpha)$ Momentum Tracks the speed of price changes for trend momentum $\text{Momentum}=\text{Close}_{t}-\text{Close}_{t-n}$ Relative Strength Index (RSI) Identifies overbought/oversold conditions using average gains and losses $\text{RSI}=100-\left(\frac{100}{1+\frac{\text{Avg Gain}}{\text{Avg Loss}}}\right)$ MACD Line Measures momentum via difference between short and long EMAs $\text{MACD}=\text{EMA}_{12}-\text{EMA}_{26}$ Williams %R Measures overbought/oversold conditions by comparing close to recent highs $\text{Williams}~\%R=\frac{\text{Highest High}_{n}-\text{Close}_{t}}{\text{Highest High}_{n}-\text{Lowest Low}_{n}}\times(-100)$ Commodity Channel Index (CCI) Identifies price deviations from a moving average for cyclical trends $\text{CCI}=\frac{\text{Price}-\text{MA}}{0.015\times\text{Mean Deviation}}$ Average Directional Index (ADX) Measures trend strength using directional movement $\text{ADX}=100\times\frac{\text{EMA}(|\text{DM}^{+}-\text{DM}^{-}|)}{\text{DM}^{+}+\text{DM}^{-}}$ Bollinger Bands Measures volatility using standard deviations around a moving average $\text{Upper Band}=\text{MA}+k\cdot\sigma$ $\text{Lower Band}=\text{MA}-k\cdot\sigma$ Stochastic Oscillator Measures momentum by comparing current close to a range of highs and lows $\%K=\frac{\text{Close}_{t}-\text{Lowest Low}_{n}}{\text{Highest High}_{n}-\text{Lowest Low}_{n}}\times 100$
Appendix B Evaluation Metrics
We evaluated the quality of the reasoning traces produced by our model using five criteria: clarity, depth, accuracy, coherence, and relevance. Each reasoning trace and its associated numeric forecast was scored on a scale from 1 (poor) to 5 (excellent). The criteria are defined as follows:

### v2

@@ 6 Conclusion
In this work, we tackled the task of verbally reasoning over financial time-series data. This task is challenging as it switches between the time-series and natural language domain for the stock price data and the reasoning step. To deal with this, we introduce our Verbal Technical Analysis (VTA) framework, which combines verbal and latent reasoning to produce interpretable time-series forecasts. The framework utilizes our Time-GRPO method to finetune the reasoning model, and conditions its forecasts on the reasoning attributes. We conducted extensive experiments and find VTA can achieve state-of-the-art forecasting accuracy while producing high-quality reasoning traces.
From the observations in this work, we propose two possible future directions. The first is to incorporate more stock characteristics into the VTA model (e.g., trend and seasonality, non-stationary behavior). The second is to improve the alignment between reasoning and forecasting, possibly through the use of more advanced techniques from image conditioning control (Zhang et al., 2023).
7 Acknowledgements
This research is supported by the Asian Institute of Digital Finance (AIDF) and the NExT Research Centre. This research is also supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant.
Appendix
Appendix A Additional Experiment Details
A.1 Model and Training Hyperparameters
This section summarizes the key hyperparameters used for training the Verbal Technical Analysis (VTA) model. All Large Language Model (LLM) components were trained using the Unsloth framework [^1] https://unsloth.ai/blog/r1-reasoning, which supports 4-bit quantization and Low-Rank Adaptation (LoRA) (Hu et al., 2022). The implementation of Group Relative Preference Optimization (GRPO) follows the principles outlined by the Hugging Face TRL library [^2] https://huggingface.co/docs/trl/main/en/grpo_trainer.
Time-Series Reasoning. The reasoning model was developed from Qwen2.5-7B-Instruct using a multi-stage training pipeline consisting of GRPO and supervised fine-tuning (SFT). The maximum sequence length was controlled via the max_seq_length parameter, and inference was performed with a temperature of 0.2.
LoRA was applied with a specific lora_rank, targeting key modules such as the attention projections (q_proj, k_proj, v_proj, o_proj) and feed-forward layers (gate_proj, up_proj, down_proj). During the initial and final GRPO training phases—used for format learning—a learning rate of $5\times 10^{-6}$ was used over two epochs. Each device processed a batch size of 4, with two gradient accumulation steps. For GRPO, four generations were produced per prompt, and the combined prompt and completion length was capped at 500 tokens. Rewards were based on adherence to the target format and accuracy of the prediction.
For SFT data generation, a rejection sampling mechanism was employed to select the top 10% Mean Squared Error (MSE) examples from a total of 100 buckets. This was followed by reasoning enhancement through SFT, using a learning rate of $2\times 10^{-4}$ over two epochs. Here, the per-device batch size was reduced to 1, while increasing the number of gradient accumulation steps to 4.
Time-Series Forecasting. The forecasting component of the VTA model is adapted from GPT-2 and uses a fixed input and output sequence length of 10 days (denoted as $T$ and $T^{\prime}$). This component was trained for 20 epochs with a learning rate of $1\times 10^{-4}$ and a batch size of 16. The architecture comprises 6 GPT-style transformer layers, as defined by the gpt_layers parameter.
LoRA was configured with a rank of 8, a scaling factor (lora_alpha) of 32, and a dropout rate of 0.1. The joint conditional training was implemented with a probability of unconditional training $p_{\text{uncond}}=0.30$, and a guidance scale of $s=0.1$. Additionally, alignment loss was incorporated using the statistical properties (i.e., minimum, maximum, and mean) of the predicted sequences.
A.2 Computational Resources
[…]
Once reasoning traces are generated, downstream forecasting runs are significantly more efficient. Training a forecasting model for a single stock with approximately 1000 training and 250 test points takes around 3 minutes on a single GPU (using $\sim$2.7 GB VRAM). A complete run over the full StockNet dataset requires approximately 4.5 hours on one GPU.
Appendix B List of Technical Indicators
Table 5: A list of the financial technical indicators used in Time-GRPO.
Indicator Description Formula
Simple Moving Average (SMA) Identifies trend by smoothing price data over a period $\text{SMA}=\frac{1}{n}\sum_{i=1}^{n}\text{Price}_{i}$
Exponential Moving Average (EMA) Measures the trend by smoothing price data with greater weight to recent prices $\text{EMA}_{t}=\text{Price}_{t}\cdot\alpha+\text{EMA}_{t-1}\cdot(1-\alpha)$
Momentum Tracks the speed of price changes for trend momentum $\text{Momentum}=\text{Close}_{t}-\text{Close}_{t-n}$
Relative Strength Index (RSI) Identifies overbought/oversold conditions using average gains and losses $\text{RSI}=100-\left(\frac{100}{1+\frac{\text{Avg Gain}}{\text{Avg Loss}}}\right)$
MACD Line Measures momentum via difference between short and long EMAs $\text{MACD}=\text{EMA}_{12}-\text{EMA}_{26}$
Williams %R Measures overbought/oversold conditions by comparing close to recent highs $\text{Williams}~\%R=\frac{\text{Highest High}_{n}-\text{Close}_{t}}{\text{Highest High}_{n}-\text{Lowest Low}_{n}}\times(-100)$
Commodity Channel Index (CCI) Identifies price deviations from a moving average for cyclical trends $\text{CCI}=\frac{\text{Price}-\text{MA}}{0.015\times\text{Mean Deviation}}$
Average Directional Index (ADX) Measures trend strength using directional movement $\text{ADX}=100\times\frac{\text{EMA}(|\text{DM}^{+}-\text{DM}^{-}|)}{\text{DM}^{+}+\text{DM}^{-}}$
Bollinger Bands Measures volatility using standard deviations around a moving average $\text{Upper Band}=\text{MA}+k\cdot\sigma$ $\text{Lower Band}=\text{MA}-k\cdot\sigma$
Stochastic Oscillator Measures momentum by comparing current close to a range of highs and lows $\%K=\frac{\text{Close}_{t}-\text{Lowest Low}_{n}}{\text{Highest High}_{n}-\text{Lowest Low}_{n}}\times 100$
Appendix C Statistical Significance of Main Results
[…]
$t=\frac{\bar{x}-\mu_{0}}{s/\sqrt{n}},$
where $\mu_{0}$ is the error of the second-best model. We reject the null hypothesis for $p<0.05$.
Table 6: One-sample, one-sided $t$-test results against the second-best model.
Metric $\bar{x}$ $s$ $t$ $p$ (one-sided)
MSE 0.06661 0.00054 4.42-4.42 $8.3\times 10^{-4}$
MAE 0.17081 0.00047 20.04-20.04 $4.5\times 10^{-9}$
Both tests yield $p\ll 0.05$, indicating that the reductions in error relative to the second-best model are statistically significant. For completeness, we report the full distribution over seeds as:
$\mathrm{MSE}=0.06661\pm 0.00054,\quad\mathrm{MAE}=0.17081\pm 0.00047\quad(\text{mean}\pm\text{std},\,n=10).$
We select the single best run (MSE = 0.06594, MAE = 0.17008) for all subsequent evaluations.
Appendix D Details on Expert Evaluation
D.1 Supplementary Information
We conducted significance tests for our expert evaluation results:
Table 7: Significance testing (paired $t$-tests). Values indicate $p$-values for pairwise comparisons.
Comparison Clarity Depth Accuracy Coherence Relevance Overall
VTA vs GPT-4.1 mini 0.0765 $9.1\times 10^{-10}$ $4.4\times 10^{-6}$ 0.0280 $1.6\times 10^{-5}$ $1.3\times 10^{-6}$
VTA vs Deepseek R1 0.0024 $1.8\times 10^{-7}$ $3.1\times 10^{-5}$ 0.0034 $1.0\times 10^{-5}$ $2.3\times 10^{-6}$
GPT-4.1 mini vs Deepseek R1 0.0555 0.0045 0.0247 0.4594 0.6854 0.8515
A statistical analysis using paired $t$-tests shows that the differences between VTA and the other two models are significant at the 5% level for all criteria except Clarity, where the base LLMs are already good at. Thus, the higher expert ratings for our model’s reasoning are statistically robust.
[…]
Experts highlighted the strengths of VTA in using a wider variety of relevant indicators and in providing conclusions that align with the explanation (e.g., “The analysis included a variety of different indicators and used them to create a coherent story”, “Interesting use and relevance of indicators”). In contrast, outputs from Deepseek-R1 and GPT-4.1 mini tended to be either vague in analysis depth or not clearly linked to the price forecast, especially in the case of Deepseek-R1.
On the other hand, experts suggested that VTA could further improve by discussing indicators in more detail and by using clearer formatting (e.g., “Should use bullet points for readability”, “Should specify explicit indicator thresholds”). While better formatting and longer explanations can be addressed in future work, precise thresholds could be more difficult to enforce since the model learns them adaptively from data.
D.2 Survey Participants
For the expert evaluation, we have surveyed domain experts with professional experience in financial market analysis. A list of their surveyed background are as follows (duplicates are removed):
J.P. Morgan, UBS, Allianz Global Investors, Evercore, Perella Weinberg Partners, Bain & Company, McKinsey & Company, L.E.K. Consulting, H.I.G. Capital, Simon Kucher, Axxion S.A., TCG Corporate Finance, BayernLB, Spinone Capital, Phillip Nova, Check24, SD Guthrie, MSMIF, Asian Institute of Digital Finance, Harvard Business School
Appendix E Additional Case Studies
Figure 9: VTA reasoning about a moderate increase (from the last price) with some pullbacks.
Figure 10: VTA reasoning about prices to be slightly lower than last price but within recent range.
Figure 11: VTA reasoning about a slight increase in price but staying within the current range.
Figure 12: VTA reasoning about a potential short-term bounce, which realized in the ground truth.
Figure 13: VTA reasoning about a slight increase, with potential for a pullback (i.e., trend reversal).
Appendix F LLM-as-a-Judge
To ensure reproducibility of the evaluation method, we further evaluated the quality of the reasoning generated by VTA using LLM-as-a-judge (Zheng et al., 2023; Gu et al., 2024). To do this, we first randomly sample 1,000 reasoning traces for evaluation. The reasoning samples are then evaluated using a stronger model as the judge, GPT-4.1 (OpenAI, 2025). The samples are judged on the same metrics on a scale of 1-5 and the average scores over all samples is reported.
Table 8: Comparison on reasoning quality across different VTA variants and reasoning LLMs.
Clarity Depth Accuracy Coherence Relevance Overall
GPT-4.1 mini 4.06 1.87 2.37 4.05 2.29 2.93
Deepseek-R1 3.95 3.02 2.97 3.98 3.25 3.43
VTA (Ours) 4.21 4.14 3.96 4.57 4.58 4.29
From Table 8, we observe that there is a clear improvement in all metrics from the inference-only reasoning models to our VTA model, showing the effectiveness of the fine-tuning process. Importantly, when comparing these LLM-as-judge results to the human experts ratings, we also find that the relative differences are highly consistent: VTA shows the largest margin over baselines in Depth, Accuracy, and Relevance, while the smallest gap is seen in Coherence and Clarity.
Appendix G Comparison in other domains
Table 9 reports the performance comparison on the Healthcare (ILI) and Energy (ETTh1) datasets.
Table 9: Performance comparison on time-series from other domains.
ILI ETTh1
MSE MAE MSE MAE
Informer 1.7372 0.8669 0.3128 0.4848
Transformer 1.6827 0.8122 0.2505 0.4360
Crossformer 1.3320 0.7270 0.1298 0.2914
TSMixer 2.1163 0.9504 0.0877 0.2375
Reformer 1.5554 0.8134 0.2422 0.4020
LightTS 2.3121 1.0612 0.0625 0.1921
DLinear 2.9160 1.3933 0.0404 0.1499
FiLM 2.6761 1.3856 0.0423 0.1543
Non-stationary 1.4328 0.8538 0.0403 0.1530
MICN 1.3431 0.8732 0.3063 0.4649
Autoformer 1.7554 1.0350 0.0736 0.2146
TimesNet 1.0795 0.6910 0.0410 0.1572
CALF 1.4442 0.7409 0.0361 0.1411
VTA (Ours) 1.4366 0.7090 0.0346 0.1374
Here, we observe that the healthcare and energy datasets do not follow the same performance patterns we saw in financial time-series, which could be attributed to different defining characteristics. For example, models that benefit from trend-seasonal decomposition on financial data do not show the same advantage here, likely because ILI and ETTh1 do not exhibit the same cyclical structure.
In these domains, especially on the ILI dataset, our VTA model also does not exhibit performance that is too far away from the time-series LLM model (CALF). It is possible that the additional reasoning traces provide limited benefit when there is not much complex signals to reason over outside of simple trend extrapolation, which was previously visualized in Figure 8.
Appendix H Financial Analysis of Portfolios
To assess the robustness of our portfolio returns, we apply standard performance attribution models widely used in finance. These models allow us to separate returns that can be explained by common risk exposures from those that reflect potential strategy-specific value.
  • Capital Asset Pricing Model (CAPM): A baseline model that relates portfolio returns to overall market returns. It provides a measure of whether the strategy delivers excess returns (alpha) after adjusting for market risk (beta).
  • Fama–French multi-factor models: Extensions of CAPM that incorporate additional risk factors, such as company size, value versus growth, profitability, investment patterns, and momentum. These factors capture well-documented drivers of returns beyond market exposure.
These models enables us to evaluate whether our results are explained by general market and factor exposures or whether they demonstrate incremental performance beyond established benchmarks.
Following industry practice, we conduct an in-depth performance attribution study of the portfolio formed from our VTA method, using both CAPM and the Fama-French 6-factor model.
H.1 The CAPM Regression Model
For CAPM, the following model was applied to the daily returns of our portfolio and the market:
$(R_{\text{VTA},t}-R_{f,t})=\alpha+\beta\cdot(R_{\text{market},t}-R_{f,t})+\epsilon_{t},$
where:
  • $(R_{\text{portfolio},t}-R_{f,t})$: The excess return of our portfolio on day $t$.
  • $(R_{\text{market},t}-R_{f,t})$: The excess return of the market benchmark on day $t$.
  • Alpha ($\alpha$): The regression intercept, which represents the portion of the portfolio’s return that is not explained by market movements.
  • Beta ($\beta$): The regression slope, which measures the portfolio’s systematic risk relative to the market.
  • Epsilon ($\epsilon_{t}$): The error term for day $t$.
[…]
We regressed our daily portfolio returns against the excess market return for each region, using the representative market indices. Below is a summary of results across the four datasets:
Table 10: CAPM Regression Results.
Dataset Market Index Used Beta ($\beta$) Annualized Alpha ($\alpha$) Alpha p-value R-squared
dowjones_30 Dow Jones Industrial (DJI) 0.6745 +3.41% 0.701 44.04%
stocknet MSCI World (ACWI) 0.6312 +10.12% 0.198 36.28%
ftse_china_a50 FTSE China A50 ETF (2822.HK) 0.3562 +53.91% 0.008 26.72%
eurostoxx_50 EURO STOXX 50 (STOXX50E) 0.2804 +15.44% 0.120 15.93%
The CAPM regression results provide a useful first diagnostic for understanding the risk-adjusted performance of our strategy. Across all four datasets, the strategy exhibits positive annualized alpha, which suggests that returns exceed those predicted by exposure to the overall market alone. However, statistical significance is only achieved in the China A50 dataset, where the alpha is both high (+53.91%) and significant ($p$ = 0.008). In other markets, while alpha remains positive, the higher $p$-values suggest weaker evidence of systematic outperformance.
Beta values in the range of 0.28 to 0.67 indicate moderate market exposure. This shows that the strategy is not entirely market-neutral, but it is also not simply tracking index movements. This aligns with our use of short-term technical forecasts rather than macro-driven positions.
The R-squared values, which range from 15.93% to 44.04%, show that a significant portion of return variation is not explained by CAPM, particularly in the European and Chinese markets. This points to a meaningful degree of idiosyncratic return generation, consistent with a model that is extracting useful trading signals beyond traditional market risk.
H.2 The Fama–French 6-Factor Model
We further evaluated performance using the Fama-French 6-factor model. This model expands on CAPM by adding five more factors: market, size, value, profitability, investment, and momentum:
$R_{p}-R_{f}=\alpha+\beta_{1}(R_{m}-R_{f})+\beta_{2}\cdot\text{SMB}+\beta_{3}\cdot\text{HML}+\beta_{4}\cdot\text{RMW}+\beta_{5}\cdot\text{CMA}+\beta_{6}\cdot\text{WML}+\epsilon$
Table 11: Fama–French 6-Factor Regression Results.
Dataset Annualized Alpha ($\alpha$) Alpha p-value R-squared
dowjones_30 +0.18% 0.985 38.77%
stocknet +9.57% 0.213 39.37%
ftse_china_a50 +73.40% 0.001 13.81%
eurostoxx_50 +13.67% 0.184 12.35%
The Fama-French 6-factor analysis gives a more granular view of performance. While positive alpha persists in all datasets, their signficance varies, ranging from 0.001 in China A50 (very significant) to 0.985 (not significant). One possible reason for the lack of significance could be due to the short period horizon of our test dataset (1 year).
The R-squared values, ranging from 12.35% to 39.37%, indicate that even after accounting for a broader set of risk factors beyond CAPM, a significant share of return variation remains unexplained, which is good. In our work, technical analysis is typically designed for short-term optimization, capturing brief momentum, reversal, or volume-based effects. Because of this short-term focus, our model is not expected to align closely with long-horizon economic models like CAPM or Fama-French, which are typically evaluated over months or quarters.
Overall, the results demonstrate that our approach does offer some complementary value to CAPM and Fama-French 6 factors to be used as an interpretable, forward-looking portfolio signal.